Find the area of the parallelogram with the given vertices. K(3, 2, 1), L(3, 4, 3), M6, 10, 3), (6, 8, 1). 10V3
Answer #1

N 161811) M 1 619,3) *1312,1 113193) ki oł tejto KN = i +6j to so area= I KLXKN | ki x Kon : Ï Ï Â | o 2 2 1 3 bol - KLXKN